ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

19. History of Mathematics

1991 MS Classification: 1, 11 Siu, Man-Keung, University of Hong Kong, Hong Kong SAR, China, Han, Qi, Institute of History of Natural Sciences, Academia Sinica, Beijing, China On the myth of an ancient Chinese theorem about primality A problem, commonly attributed to the ancient Chinese, is about the primality of n and the congruence between 2^{n} and 2 modulo n. This myth seems to have been caused by a note published by the British missionary Alexander Wylie in 1869 in a journal in Hong Kong, which reported on a theorem of Li Shanlan. We will examine this story in more details.